OpenStudy (anonymous):

Solve the quadratic equation. 2x2 + 35 = 17x

OpenStudy (anonymous):

answer choices A) 7/2and 5 B) -7\2and 5 C) -7\2and -5 D) 2/7and 1/5

OpenStudy (jiteshmeghwal9):

2x^2+35-17x=0 now divide the equation by the coefficient of x^2 \(\LARGE{{2x^2+35-17x \over 2}=\dfrac{0}{2}}\) \(\LARGE{x^2+\dfrac{35}{2}-\dfrac{17x}{2}=0}\) \(\LARGE{x^2+({-17/2}) \space x=-35/2}\)

OpenStudy (hba):

There are three methods we can implement. 1) Using the quad formula 2)breaking the middle term 3)Completing the square method.

OpenStudy (hba):

For Using The Quad formula you would have to compare your equation with the standard equation. Which is, \[\huge\ ax^2+bx+c=0\] And determine a,b and c.

OpenStudy (jiteshmeghwal9):

\[x^2-\dfrac{17}{2}x +\left( 17 \over 4 \right)^2={-35 \over 2}+\left( 17 \over 4 \right)^2\]

OpenStudy (hba):

Then Using the Quad formula, \[\huge\ x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] @jiteshmeghwal9 has showed completing the square method.

OpenStudy (jiteshmeghwal9):

\[(x+\dfrac{17}{4})^2={-35 \over 2}+\dfrac{289}{16}\]

OpenStudy (jiteshmeghwal9):

Now can u do it further @Pookiieluvzyhu ??

OpenStudy (hba):

Looks like @jazy is gonna take a day lol :P

OpenStudy (anonymous):

yea yea..kk thanks...let me read over what yall said.it was alot lollz.. thanks @jiteshmeghwal9 @hba @jazy

OpenStudy (jiteshmeghwal9):

yw :)

OpenStudy (anonymous):

Well, you could use the Quad formula. Quadratic equation - \[ax^2 + bx + c\]You have...\[2x^2 - 17x + 35\] a = 2 b = -17 c = 35 Quadratic Formula.\[\frac{ -b \pm \sqrt{b^2 - 4ac} }{2a }\]Plug in the values.\[\frac{ 17 \pm \sqrt{49 - 4(2)(35)}}{ 2(2) } \]Solve.

OpenStudy (hba):

Welcome :)

OpenStudy (anonymous):

Wow that took a longgggg time to post O_O

OpenStudy (anonymous):

Welcome if it helps lol. (: